A chemist prepares a solution of iron(III) bromide FeBr3 by measuring out 57.3g of iron(III) bromide into a 50.mL volumetric flask and filling the flask to the mark with water.

Calculate the concentration in mol/L of the chemist's iron(III) bromide solution. Be sure your answer has the correct number of significant digits.
Answer #1

mass of Iron till) bromide =57039 (Fe Brg) Molar mass of fe Bory = 56+3x80 = 56+240 = 290g/mole. no. of moles of Fe Brz = giv